OpenStudy (anonymous):

what is 3 - 2(8 + x) -4x simplified?

OpenStudy (anonymous):

-13-2x

OpenStudy (anonymous):

3-2(8+x)-4x = 3-16-2x-4x = -13-6x = -6x-13